OpenStudy (anonymous):

In order for g(x) to have no points of discontinuity a must =? g(x)={ax x<2 and 9 x>or equal to 2

OpenStudy (inkyvoyd):

for x =2 both piecewise functions must be equal.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we want \(ax=9\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

When \(x=2\) or something

OpenStudy (inkyvoyd):

what @wio said.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve for \(a\).

OpenStudy (anonymous):

a=7

OpenStudy (anonymous):

7(2)=9? I want to believe

OpenStudy (inkyvoyd):

No, a*2=9 lol wio

OpenStudy (anonymous):

I am confused becuase 9-2=7 so then shouldn't a=7

OpenStudy (inkyvoyd):

No, you are solving the equation 2a=9, not 2+a=9.

OpenStudy (anonymous):

so the would a=4.5

OpenStudy (inkyvoyd):

Yes...

OpenStudy (anonymous):

ok thanks